Solve the equation
0.32(24)+0.05x=0.15(18+x)

Answer by stanbon(75887)

Solve the equation
0.32(24)+0.05x=0.15(18+x)
Multiply thru by 100 to get rid of the decimals:
32*24 + 5x = 15*18 + 15x
10x = 32*24 - 15*18
x = 49.8
===============
Cheers,
Stan H.